# Multi-Fidelity Bayesian Optimization for Nash Equilibria with Black-Box Utilities

### v1

@@ Abstract
Modern open and softwarized systems – such as O-RAN telecom networks and cloud computing platforms – host independently developed applications with distinct, and potentially conflicting, objectives. Coordinating the behavior of such applications to ensure stable system operation poses significant challenges, especially when each application’s utility is accessible only via costly, black-box evaluations. In this paper, we consider a centralized optimization framework in which a system controller suggests joint configurations to multiple strategic players, representing different applications, with the goal of aligning their incentives toward a stable outcome. To model this interaction, we formulate a Stackelberg game in which the central optimizer lacks access to analytical utility functions and instead must learn them through sequential, multi-fidelity evaluations. To address this challenge, we propose MF-UCB-PNE, a novel multi-fidelity Bayesian optimization strategy that leverages a budget-constrained sampling process to approximate pure Nash equilibrium (PNE) solutions. MF-UCB-PNE systematically balances exploration across low-cost approximations with high-fidelity exploitation steps, enabling efficient convergence to incentive-compatible configurations. We provide theoretical and empirical insights into the trade-offs between query cost and equilibrium accuracy, demonstrating the effectiveness of MF-UCB-PNE in identifying effective equilibrium solutions under limited cost budgets.
Index Terms:
Bayesian optimization, Nash equilibria, multi-fidelity simulation, Gaussian bandits
@@ I-A Context and Motivation
In open, softwarized, systems, different applications are run simultaneously on shared physical resources while pursuing separate, and potentially conflicting, objectives. As illustrated in Fig. 1, an important example is given by open telecom network architectures such as O-RAN [1], in which apps implementing network functionalities are designed to optimize distinct objectives by controlling physical and virtual parameters such as transmission powers and computational resources [2, 3]. Other examples include cloud computing environments involving tenants with diverse service level agreements [4], and scheduling policy selection for service slices [5].
@@
The potentially conflicting nature of the objectives of different apps may yield instabilities and undesirable behaviors [2]. For example, in the setting of Fig. 1, one app aims at controlling energy consumption by adaptively shutting down a transmitter, while another attempts to maximize the communication throughput by increasing energy consumption. This may yield reduced performance levels when both apps are deployed simultaneously.
Motivated by the management of conflicts between independent apps in complex systems like telecom networks, this paper studies the setting depicted in Fig. 2. In it, a central optimizer suggests actions to different players, representing the apps in the motivating examples above. The goal for the optimizer is to ensure that the players have no, or limited, incentives to deviate from the suggested actions. Such set of actions is thus compatible with the individual incentives of the players. This guarantees a stable operation of the system as the players adhere to the recommendations of the central optimizer.
This type of setting may be modeled as a hierarchical, or Stackelberg, game, in which the actions of competing players are, at least partially, controlled by a common optimizer [6, 7, 8]. However, in conventional hierarchical games, the utilities of the players are assumed to be known to the centralized optimizer. This is not the case where the optimizer needs to control black-box apps designed by third parties [9]. For instance, in the O-RAN architectures, apps are ideally designed and provided by a variety of vendors, diversifying the supply chain for telecom networks [10]. Therefore, in such settings, the utility functions must be modeled as black-box functions.
@@ I-B Related Work
Centralized Nash equilibrium evaluation with unknown utility functions: When analytical expressions of the utility functions are unknown, the centralized evaluation of a NE must rely on observations of the utilities for involved players. Reference [27] characterized the query complexity of finding a NE in various game classes; while references [28, 29] introduced a BO-based optimizer that can obtain approximate NE with formal regret guarantees. However, these prior works do not account for the possibility to leverage multi-fidelity observations.
@@ I-C Main Contributions
Motivated by modern softwarized systems such as open telecom networks (see Fig. 1), this paper addresses the setting in Fig. 2, in which a centralized optimizer recommends a joint configuration for the actions of a number of players so that the players have limited incentives to deviate from the suggested actions. To this end, we assume that the centralized optimizer has access to multi-fidelity estimates of the expensive-to-evaluate black-box utility functions of the players with ordered query costs.
In this context, we introduce MF-UCB-PNE, a novel MFBO policy that, under a limited query budget, aims at approaching a solution that approximately satisfies pure Nash equilibrium (PNE) conditions. MF-UCB-PNE operates in episodes that encompass exploration – to acquire information about the utility functions from low fidelity observations – as well as exploitation – to evaluate the quality of the current solutions using maximum-fidelity observations.
The main contributions are as follows:
@@
  • We provide an analysis of the regret of the proposed MF-UCB-PNE policy with respect to the utility values attainable with optimal approximate PNE solutions. The theoretical results show that MF-UCB-PNE ensures asymptotic no-regret performance given sufficiently large query budget.
@@ II-A Setting
where $\mathbf{x}_{-n}$ represents the actions performed by all players except for player $n$. As explained next, in order to ensure the stability of the action profile $\mathbf{x}$, the central agent aims at selecting a pure Nash equilibrium (PNE) of the strategic game defined by the utilities $\{u_{n}(\mathbf{x})\}_{n\in\mathcal{N}}$.
@@ II-B Optimization
The centralized optimizer produces a series of action profiles $\mathbf{x}_{t}=[\mathbf{x}_{1,t},...,\mathbf{x}_{N,t}]$ at each time $t=1,...,T$ for all $N$ players. For each choice $\mathbf{x}_{t}$, the optimizer receives a noisy estimate of the utilities $\{u_{n}(\mathbf{x}_{t})\}_{n=1}^{N}$. Specifically, we adopt a multi-fidelity model, whereby the utilities can be measured at one of $M$ fidelity levels =$\mathcal{M}=\{1,...,M\}$. Level $m=1$ corresponds to the lowest-fidelity measurement, while $m=M$ to the highest-fidelity level. Different fidelity levels may correspond to simulators with varying levels of accuracy for physical experiments, or to varying precision levels of measuring instruments. Accordingly, the central optimizer selects a fidelity level vector $\mathbf{m}_{t}=[m_{1,t},...,m_{N,t}]$ at each time $t$, along with the action profile $\mathbf{x}_{t}$, to form a decision pair $(\mathbf{x}_{t},\mathbf{m}_{t})$.
@@
$\lambda^{(1)}\leq\lambda^{(2)}\leq...\leq\lambda^{(M)}.$ (9)
The total query cost is constrained by a pre-determined total query budget $\Lambda$ as
@@
Given the pre-determined query budget $\Lambda$, our goal is to design a selection strategy $\pi$ mapping the collected observations $\mathcal{D}_{t-1}=\cup_{n\in\mathcal{N}}\mathcal{D}_{n,t-1}$ to the decision pair $(\mathbf{x}_{t},\mathbf{m}_{t})$ at all time steps $t=1,...,T$, such that the optimizer approaches an $\epsilon^{*}$-PNE solution in the set (4). To this end, as in the single-objective framework studied in [19], we define a proxy optimization objective in terms of a loss function that accounts for the dissatisfaction level associated with the solutions explored by the optimizer as well as for the cost of exploration. Specifically, we distinguish between exploration and evaluation optimization rounds:
  • Evaluation round: In an evaluation round $t$, the fidelity levels $\mathbf{m}_{t}$ are set at the maximum level $M$ for all players, i.e., $\mathbf{m}_{t}=M\cdot\mathbf{1}$, where $\mathbf{1}$ is the all one vector. This allows the optimizer to estimate the actual utilities. At the evaluation round $t$, the accrued loss is the dissatisfaction level $\max_{n\in\mathcal{N}}f_{n}(\mathbf{x}_{t})$ associated with the selected action profile $\mathbf{x}_{t}$.
  • Exploration round: In an exploration round $t$, there exists at least one fidelity level $m_{n}$ smaller than $M$ for some player $n\in\mathcal{N}$. In this case, the loss equals a constant $C$ representing the cost of exploration at lower fidelity levels. To avoid trivial solutions, we assume that this cost is larger than the largest dissatisfaction level, i.e.,
    $C\geq\max\limits_{\begin{subarray}{c}n\in\mathcal{N}\\ \mathbf{x}\in\mathcal{X}\end{subarray}}f_{n}(\mathbf{x}).$ (11)
Overall, given a selection $(\mathbf{x},\mathbf{m})$, the loss accrued by the optimizer is defined as
$\ell(\mathbf{x},\mathbf{m})=\begin{cases}\max\limits_{n\in\mathcal{N}}f_{n}(\mathbf{x})&\text{if}\,\,\mathbf{m}=M\cdot\mathbf{1}\quad\text{(evaluation)}\\ C&\text{otherwise}\quad\text{(exploration).}\end{cases}$ (12)
@@
$\min\limits_{\pi,T}\sum_{t=1}^{T}\ell(\pi(\mathcal{D}_{t-1}))$ (13a)
[…]
under the budget cost (10), where the policy $\pi$ return
@@
at each time $t=1,...,T.$ Note that the maximum number of time steps $T$ in (13) is not fixed as a priori, as it can be optimized under the budget constraint (13b).
An idealized central optimizer would know at least one action profile $\mathbf{x}^{*}$ in the target $\epsilon^{*}$-PNE set $\mathcal{X}^{(\epsilon^{*})}$. Such an optimizer could hence choose the action profile $\mathbf{x}^{*}\in\mathcal{X}^{(\epsilon^{*})}$ for all time steps allowed by the budget, obtaining a loss value (13a) equal to $\Lambda\epsilon^{*}/(N\lambda^{(M)})$. In fact, the maximum number of times that action profile $\mathbf{x}^{*}$ can be selected is $\Lambda/(N\lambda^{(M)})$, since the cost of such an action profile is $N\lambda^{(M)}$. Therefore, the quality of a policy $\pi$ designed to target problem (13) can be measured by the regret
$R(\Lambda)=\sum_{t=1}^{T}\ell(\pi(\mathcal{D}_{t-1}))-\frac{\Lambda\epsilon^{*}}{N\lambda^{(M)}}.$ (15)
@@ III-A Overview of MF-UCB-PNE
$\Lambda_{j}=\Lambda_{\mathcal{L}_{j}}+N\lambda^{(M)},$ (17)
@@
and $N\lambda^{(M)}$ being the cost of the evaluation phase.
@@
2 while $\Lambda^{\text{tot}}_{j}\geq N\lambda^{(M)}$ do
3       Run MF-UCB-PNE-Episode$(\Lambda^{\text{tot}}_{j},\{\mathcal{D}_{n,j,0}\}_{n\in\mathcal{N}})$ (Algorithm 2) to obtain the sequence of decision pairs $\mathcal{L}_{j}$, observations $\{\mathbf{y}_{n,j,t_{j}}\}_{n\in\mathcal{N}}$, accumulated cost $\Lambda_{\mathcal{L}_{j}}$ and updated dataset $\{\mathcal{D}_{n,j,t_{j}}\}_{n\in\mathcal{N}}$
@@ 7       Select the exploring action profile $\mathbf{x}^{(n_{j})}$ via (35)
11       Calculate the remaining budget $\Lambda^{\text{tot}}_{j+1}=\Lambda^{\text{tot}}_{j}-\Lambda_{\mathcal{L}_{j}}-N\lambda^{(M)}$
@@ III-B Multiple-Output Gaussian Process
where $10<\rho^{(m)}<$ 1 is a correlation coefficient, and we have $u_{n}^{(M)}(\mathbf{x})=u_{n}(\mathbf{x})\sim\mathcal{GP}(0,\kappa(\mathbf{x},\mathbf{x}^{\prime}))$, and $q^{(m)}_{n}(\mathbf{x})\sim\mathcal{GP}(0,\kappa^{(m)}(\mathbf{x},\mathbf{x}^{\prime}))$, where the kernel functions are often instantiated as the radial basis function (RBF) kernels
$\kappa(\mathbf{x},\mathbf{x}^{\prime})=\exp(-h||\mathbf{x}-\mathbf{x}^{\prime}||^{2}),$ (20)
$\kappa^{(m)}(\mathbf{x},\mathbf{x}^{\prime})=\exp(-\zeta^{(m)}||\mathbf{x}-\mathbf{x}^{\prime}||^{2})$ (21)
@@
Overall, the MOGP mean $\mu^{(m)}_{n,t}(\mathbf{x})$ in (22) provides an estimated value for the approximation $u_{n}^{(m)}(\mathbf{x})$ based on the collected dataset (8), while the variance $[\sigma_{n,t}^{(m)}(\mathbf{x})]^{2}$ quantifies the corresponding uncertainty of the estimate.
@@ III-C Exploration Phase
With this background in place, we can now define the two phases implemented by MF-UCB-PNE in each episode $\mathcal{E}_{j}$ (see Fig. 3). In the exploration phase, we adopt an information-theoretic criterion for selecting the sequence of actions. To elaborate, denote as $\mathcal{D}_{n,j,\tau-1}$ the data collected up to time $\tau-1$ in episode $j$. By defined next, at the beginning of time $\tau\geq 1$ in episode $\mathcal{E}_{j}$, the optimizer’s policy $\pi$ selects the next decision pair $(\mathbf{x}_{j,\tau},\mathbf{m}_{j,\tau})$ so as to maximize the ratio between the overall information obtained about the true value of the utility functions $\{u_{n}(\mathbf{x})\}_{n\in\mathcal{N}}$ and the cost of the selection $(\mathbf{x}_{j,\tau},\mathbf{m}_{j,\tau})$.
@@
where the GP posterior variance $[\sigma_{n,j,\tau-1}^{(m_{n,j,\tau})}(\mathbf{x}_{j,\tau})]^{2}$ is given in (23). The mutual information (24) measures the uncertainty reduction about utility $u_{n}(\mathbf{x}_{j,\tau})$ that can be obtained by choosing the pair $(\mathbf{x}_{j,\tau},\mathbf{m}_{j,\tau})$ at time step $\tau$ in episode $j$.
@@
    $\Lambda^{\text{tot}}_{j}<N(\lambda^{(1)}+\lambda^{(M)}),$ (27)
@@
[…]
    The constraint (28b) ensures that the remaining query budget after the current query is sufficient for at least an evaluation step at the maximum fidelity level, which has cost $N\lambda^{(M)}$. Having solved problem (28), the following options may be realized.
@@
      Small accumulated information gain-cost ratio: If the information gain-cost ratio accumulated on selected sequence $\mathcal{L}_{j,\tau}=\{(\mathbf{x}_{j,1},\mathbf{m}_{j,1}),...,\\ (\mathbf{x}_{j,\tau},\mathbf{m}_{j,\tau})\}$ in current episode $j$ satisfies the inequality

### v2

@@ Abstract
Modern open and softwarized systems – such as O-RAN telecom networks and cloud computing platforms – host independently developed applications with distinct, and potentially conflicting, objectives. Coordinating the behavior of such applications to ensure stable system operation poses significant challenges, especially when each application’s utility is accessible only via costly, black-box evaluations. In this paper, we consider a centralized optimization framework in which a system controller suggests joint configurations to multiple strategic players, representing different applications, with the goal of aligning their incentives toward a stable outcome. This interaction is modeled as a learned optimization with an equilibrium constraint in which the central optimizer learns the utility functions through sequential, multi-fidelity evaluations with the goal of identifying a pure Nash equilibrium (PNE). To address this challenge, we propose MF-UCB-PNE, a novel multi-fidelity Bayesian optimization strategy that leverages a budget-constrained sampling process to approximate PNE solutions. MF-UCB-PNE systematically balances exploration across low-cost approximations with high-fidelity exploitation steps, enabling efficient convergence to incentive-compatible configurations. We provide theoretical and empirical insights into the trade-offs between query cost and equilibrium accuracy, demonstrating the effectiveness of MF-UCB-PNE in identifying effective equilibrium solutions under limited cost budgets.
@@ I-A Context and Motivation
In open, softwarized, systems, different applications are run simultaneously on shared physical resources while pursuing separate, and potentially conflicting, objectives. As illustrated in Fig. 1, an important example is given by open telecom network architectures such as O-RAN [33], in which apps implementing network functionalities are designed to optimize distinct objectives by controlling physical and virtual parameters such as transmission powers and computational resources [13, 48]. Examples include cloud computing environments involving tenants with diverse service level agreements [29], and scheduling policy selection for service slices [32].
@@
The potentially conflicting nature of the objectives of different apps may yield instabilities and undesirable behaviors [13]. For example, in the setting of Fig. 1, one app aims at controlling energy consumption by adaptively shutting down a transmitter, while another attempts to maximize the communication throughput by increasing energy consumption. This may yield reduced performance levels when both apps are deployed simultaneously. Other examples include scheduling apps for the medium access control (MAC) layer, with different apps focusing on the throughput optimization of distinct users [7].
Motivated by the management of conflicts between independent apps in complex systems like telecom networks, this paper studies the setting depicted in Fig. 2. In it, a central optimizer suggests actions to different players, representing the apps in the motivating examples above. The goal for the optimizer is to ensure that the players have no, or limited, incentives to deviate from the suggested actions. Such set of actions is thus compatible with the individual incentives of the players. This ensures a stable operation of the system as the players adhere to the suggestions of the central optimizer.
This type of setting may be modeled as a learned optimization problem with equilibrium constraints, in which the actions of competing players are, at least partially, controlled by a common optimizer [30, 25, 46]. In conventional optimization with equilibrium constraints, the utilities of the players are assumed to be known to the centralized optimizer [27]. This is not the case where the optimizer needs to control black-box apps designed by third parties [47]. For instance, in the O-RAN architectures, apps are ideally designed and provided by a variety of vendors, diversifying the supply chain for telecom networks [1]. Therefore, in such settings, the utility functions must be modeled as black-box functions.
@@ I-B Related Work
This work lies at the intersection of several research domains, notably black-box optimization, under both single-fidelity [12, 43, 39, 18, 11] and multi-fidelity [38, 50, 19] data acquisition settings, and the evaluation of Nash equilibria, considering both known [45, 4, 9, 35, 8] and unknown [10, 40, 15] utility functions. We then discuss each of these research areas in turn.
@@
Centralized Nash equilibrium evaluation with unknown utility functions: When analytical expressions of the utility functions are unknown, the centralized evaluation of a NE must rely on observations of the utilities for involved players. Reference [10] characterized the query complexity of finding a NE in various game classes; while references [40, 15] introduced BO-based optimizers that can approximate NE with formal regret guarantees. However, prior works do not account for the possibility to leverage multi-fidelity observations.
@@ I-C Main Contributions
Motivated by modern softwarized systems such as open telecom networks (see Fig. 1), this paper addresses the setting in Fig. 2, in which a centralized optimizer recommends a joint configuration for the actions of a number of players so that the players have limited incentives to deviate from the suggested actions. To this end, we assume that the centralized optimizer uses multi-fidelity estimates of the expensive-to-evaluate black-box utility functions of the players with ordered query costs.
In this context, we introduce MF-UCB-PNE, a novel MFBO policy that, under a limited query budget, aims at approaching a solution that approximately satisfies pure Nash equilibrium (PNE) conditions. MF-UCB-PNE operates in episodes that encompass exploration – to acquire information about the utility functions from low fidelity observations – as well as exploitation – to evaluate the current solutions using maximum-fidelity observations. The main contributions are as follows:
@@
  • We provide the regret analysis of the proposed MF-UCB-PNE policy with respect to the utility values attainable with optimal approximate PNE solutions. The theoretical results show that MF-UCB-PNE ensures asymptotic no-regret performance given sufficiently large query budget.
@@ II-A Setting
where $\mathbf{x}_{-n}$ represents the actions performed by all players except for player $n$.
As explained next, in order to ensure the stability of the action profile $\mathbf{x}$, the central agent aims at selecting a pure Nash equilibrium (PNE) of the strategic game defined by the utilities $\{u_{n}(\mathbf{x})\}_{n\in\mathcal{N}}$. This way, the operation of the central controller can be formulated as a form of optimization with equilibrium constraints [30, 25, 46], where the equilibrium requirement pertains to the game played among the $N$ players based on the utilities in (1).
@@ II-B Optimization
The centralized optimizer produces a series of action profiles $\mathbf{x}_{t}=[\mathbf{x}_{1,t},...,\mathbf{x}_{N,t}]$ at each time t=1,,Tt=1,...,T for all $N$ players. For each choice $\mathbf{x}_{t}$, the optimizer receives a noisy estimate of the utilities $\{u_{n}(\mathbf{x}_{t})\}_{n=1}^{N}$. Specifically, we adopt a multi-fidelity model, whereby the utilities can be measured at one of $M$ fidelity levels $\mathcal{M}=\{1,...,M\}$. Level $m=1$ corresponds to the lowest-fidelity measurement, while $m=M$ to the highest-fidelity level. Different fidelity levels may correspond to simulators with varying levels of accuracy for physical experiments, or to varying precision levels of measuring instruments. Accordingly, the central optimizer selects a fidelity level vector $\mathbf{m}_{t}=[m_{1,t},...,m_{N,t}]$ at each time $t$, along with the action profile $\mathbf{x}_{t}$, to form a decision pair $(\mathbf{x}_{t},\mathbf{m}_{t})$.
@@
$\lambda^{(1)}\leq\lambda^{(2)}\leq...\leq\lambda^{(M)}=1,$ (9)
where we have set the maximum cost $\lambda^{(M)}=1$ without loss of generality. The total query cost is constrained by a pre-determined total query budget $\Lambda$ as
@@
Given the pre-determined query budget $\Lambda$, our goal is to design a selection strategy $\pi$ mapping the collected observations $\mathcal{D}_{t-1}=\cup_{n\in\mathcal{N}}\mathcal{D}_{n,t-1}$ to the decision pair $(\mathbf{x}_{t},\mathbf{m}_{t})$ at all time steps t=1,,Tt=1,...,T, such that the optimizer approaches an $\epsilon^{*}$-PNE solution in the set (4).
To this end, as in the single-objective framework studied in [38], we define a proxy optimization objective in terms of a reward function that accounts for the dissatisfaction level associated with the solutions evaluated by the optimizer at the maximum fidelity, as well as for the cost of exploration at lower fidelity levels. To proceed, fix a constant $C$ to be any upper bound on the maximum value of the dissatisfaction:
$C\geq\max\limits_{n\in\mathcal{N},\mathbf{x}\in\mathcal{X}}f_{n}(\mathbf{x}).$ (11)
The reward is then defined by distinguishing between exploration and evaluation rounds:
  • Evaluation round: In an evaluation round $t$, the fidelity levels $\mathbf{m}_{t}$ are set at the maximum level $M$ for all players, i.e., $\mathbf{m}_{t}=M\cdot\mathbf{1}$, where $\mathbf{1}$ is the all-one vector. This allows the optimizer to estimate the actual utilities of all players. At any evaluation round $t$, the accrued reward is the normalized dissatisfaction improvement
    $r(\mathbf{x}_{t})=\Big[C-\max_{n\in\mathcal{N}}f_{n}(\mathbf{x}_{t})\Big]/C$ (12)
    associated with the selected action profile $\mathbf{x}_{t}$. By (11), the reward (12) is non-negative. In particular, if (12) holds with equality, it ranges from 0 (maximum dissatisfaction) to $1-\epsilon^{*}/C$ ($\epsilon^{*}$-PNE).
  • Exploration round: In an exploration round $t$, there exists at least one fidelity level $m_{n}$ smaller than $M$ for some player $n\in\mathcal{N}$. In this case, following [38], we set the reward to 0. This way, observations at lower fidelity levels are leveraged purely for exploration, with the goal of enhancing future rewards (12) to be obtained in subsequent evaluation rounds.
Overall, given a selection $(\mathbf{x},\mathbf{m})$, the reward accrued by the optimizer is defined as
$r(\mathbf{x},\mathbf{m})=\begin{cases}\frac{C-\max_{n\in\mathcal{N}}f_{n}(\mathbf{x})}{C}&\text{if}\,\,\mathbf{m}=M\cdot\mathbf{1}\quad\text{(evaluation)}\\ 0&\text{otherwise}\quad\text{(exploration).}\end{cases}$ (13)
@@
$\max\limits_{\pi,T}\sum_{t=1}^{T}r(\pi(\mathcal{D}_{t-1}))$ (14a)
$\text{s.t.}\quad\sum_{n=1}^{N}\sum_{t=1}^{T}\lambda^{(m_{n,t})}\leq\Lambda$ (14b)
under the budget cost (10), where the policy $\pi$ return the actions
@@
at each time t=1,,T.t=1,...,T.
An idealized central optimizer would know at least one action profile $\mathbf{x}^{*}$ in the target $\epsilon^{*}$-PNE set $\mathcal{X}^{(\epsilon^{*})}$. Such an optimizer could hence avoid exploration steps, and choose to evaluate directly the action profile $\mathbf{x}^{*}\in\mathcal{X}^{(\epsilon^{*})}$ for all time steps allowed by the budget, obtaining a reward value (14a) equal to $\Lambda(1-\epsilon^{*}/C)/N$. Note, in fact, that the maximum number of evaluation steps is $\Lambda/N$, since the cost of each evaluation step equals $N$. Therefore, the quality of a policy $\pi$ designed to target problem (14) can be measured by the regret
$R(\Lambda)=\frac{\Lambda}{N}\Big(1-\frac{\epsilon^{*}}{C}\Big)-\sum_{t=1}^{T}r(\pi(\mathcal{D}_{t-1})).$ (16)
@@ III-A Overview of MF-UCB-PNE
$\Lambda_{j}=\Lambda_{\mathcal{L}_{j}}+N,$ (18)
@@
and $N$ being the cost of the evaluation phase.
@@
2 while $\Lambda^{\text{tot}}_{j}\geq N$ do
[…]
@@ 7   Select the exploring action profile $\mathbf{x}^{(n_{j})}$ via (36)
11   Calculate the remaining budget $\Lambda^{\text{tot}}_{j+1}=\Lambda^{\text{tot}}_{j}-\Lambda_{\mathcal{L}_{j}}-N$
@@ III-B Multiple-Output Gaussian Process
where $0<\rho^{(m)}<1$ is a correlation coefficient, and we have $u_{n}^{(M)}(\mathbf{x})=u_{n}(\mathbf{x})\sim\mathcal{GP}(0,\kappa(\mathbf{x},\mathbf{x}^{\prime}))$, and $q^{(m)}_{n}(\mathbf{x})\sim\mathcal{GP}(0,\kappa^{(m)}(\mathbf{x},\mathbf{x}^{\prime}))$, where the kernel functions are often instantiated as the radial basis function (RBF) kernels
[…]
@@
Overall, the MOGP mean $\mu^{(m)}_{n,t}(\mathbf{x})$ in (23) provides an estimated value for the approximation $u_{n}^{(m)}(\mathbf{x})$ based on the collected dataset (8), while the variance $\sigma_{n,t}^{(m)}(\mathbf{x})]^{2}$ quantifies the corresponding uncertainty of the estimate.
@@ III-C Exploration Phase
With this background in place, we can now define the two phases implemented by MF-UCB-PNE in each episode $\mathcal{E}_{j}$ (see Fig. 3). In the exploration phase, we adopt an information-theoretic criterion for selecting the sequence of actions. To elaborate, denote as $\mathcal{D}_{n,j,\tau-1}$ the data collected up to time $\tau-1$ in episode $j$. By defined next, at the beginning of time $\tau\geq 1$ in episode $\mathcal{E}_{j}$, the optimizer’s policy $\pi$ selects the next pair $(\mathbf{x}_{j,\tau},\mathbf{m}_{j,\tau})$ to maximize the ratio of the overall information obtained about the true value of the utility functions $\{u_{n}(\mathbf{x})\}_{n\in\mathcal{N}}$ and the cost of the selection $(\mathbf{x}_{j,\tau},\mathbf{m}_{j,\tau})$.
@@
where the GP posterior variance $\sigma_{n,j,\tau-1}^{(m_{n,j,\tau})}(\mathbf{x}_{j,\tau})]^{2}$ is given in (24). The mutual information (25) measures the uncertainty reduction about utility $u_{n}(\mathbf{x}_{j,\tau})$ that can be obtained by choosing the pair $(\mathbf{x}_{j,\tau},\mathbf{m}_{j,\tau})$ at time step $\tau$ in episode $j$.
@@
    $\Lambda^{\text{tot}}_{j}<N(\lambda^{(1)}+1),$ (28)
@@
    $\text{s.t.}\quad\sum_{n=1}^{N}\lambda^{(m_{n})}\leq\Lambda^{\text{tot}}_{j,\tau}-N.$ (29b)
    The constraint (29b) ensures that the remaining query budget after the current query is sufficient for at least an evaluation step at the maximum fidelity level, which has cost $N$. Having solved problem (29), the following options may be realized.
@@
      Small accumulated information gain-cost ratio: If the information gain-cost ratio accumulated on selected sequence $\mathcal{L}_{j,\tau}=\{(\mathbf{x}_{j,1},\mathbf{m}_{j,1}),...,\\ (\mathbf{x}_{j,\tau},\mathbf{m}_{j,\tau})\}$ in episode $j$ satisfies the inequality
@@
III-D Evaluation Phase
$\hskip 14.22636pt+\beta_{n,j}\sigma^{(M)}_{n,j,t_{j}}(\mathbf{x})],$ (32)